# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ III-C2 All-optical nonlinearities
Without electrical components, the all-optical approach implements specific activation functions via the nonlinear response of materials or devices to optical signals. Here, the optical signal may serve either as the input or modulate other optical signals. Several ONNs have been demonstrated using semiconductor optical amplifiers (SOAs) [130], saturable absorber [131, 132], and techniques exhibiting excitable behavior [133]. All-optical nonlinearities can also be achieved by PCM, such as the all-optical neuron developed by Feldmann et al. In this work, PCM, functioning as waveguide cladding, is used to modulate pre-synaptic input and govern the resonance state of MRR, thereby controlling the generation of output spikes (Fig. 8.c) [34]. Beyond the issues of activation function applicability previously discussed, the implementation of all-optical neurons faces several other challenges. Firstly, even though extra electrical devices are not required, the substantial power required to excite nonlinearities does not confer a power consumption benefit compared to the O-E-O approaches. Additionally, the fabrication and integration of optical components for all-optical nonlinearity, such as PCM, excitable lasers, and optical amplifiers, pose challenges as well. Lastly, since both the control signal and post-synaptic output are optical signals, a potential issue is distinguishing the response from the control signals or bias.
@@ III-D Optical Neural Network Training
As an analog computing platform, the PICs for AI applications are inherently susceptible to various non-ideal conditions, such as environmental changes, process variation, and limited control precision. These factors can lead to the decrease in inference accuracy and potentially degrade the signal-to-noise ratio (SNR) of computations. To mitigate the decrease in accuracy, it’s crucial to train the optical neural network with careful consideration of non-idealities that may occur during inference, making it ”circuit-aware” to improve its noise resilience. The current approaches to ONN training can be classified into two categories: hardware-aware ex-situ training, which involves training with the help of digital computers, and on-chip in-situ training.
@@ III-D1 Hardware-aware ex-situ Training
Hardware-aware ex-situ training offloads the training process of ONN to digital computers and utilizes various circuit-aware training techniques to capture circuit behavior as precisely as possible during training. One commonly used technique is noise-aware training [134, 135], as depicted in Fig. 9.a. This approach involves modeling the behavior of PIC while considering various non-ideal effects. Subsequently, the PIC model is injected during training to reduce the gap between training and real inference. Gaussian noise is commonly used to model various noise sources in photonic systems. For instance, shot noise and thermal noise are modeled using a Gaussian distribution by measuring on-chip photonic multiplication results and fitting their distribution [120]. In addition to injecting non-idealities, some work explicitly models the transfer matrix of photonic neurons and embeds them in the forward computation pass during training. For example, the transfer matrix of photonic neurons [118, 119] is modeled and injected during training. This injection is crucial as it introduces a unique nonlinearity term that ONNs need to be aware of in order to accurately capture the behavior of the photonic neurons. Moreover, in Ref [114], the noisy transfer matrix is derived and explicitly injected in the training under some noise assumptions. However, this approach encounters two main challenges. Firstly, accurately representing all on-chip non-idealities poses a significant difficulty, and environmental fluctuations can further affect inference accuracy. Second, the computational overhead required to model the physical system accurately can be exceedingly high or even prohibitive, making the training very slow and costly.
Aware of these challenges, some works advocate for training optical neural networks directly with non-ideal physical responses, called physical network training, as shown in Fig. 9.b. The forward pass is executed on PICs and the loss signal $\mathcal{L}$ is obtained by comparing the physical and intended outputs. However, the challenge arises during the backward pass due to the undifferentiable nature of the physical response, making backpropagation difficult. To address this challenge, the idea of adopting a differentiable digital PIC surrogate model has been proposed [81, 136, 137]. With the differentiable digital model, the gradient of the loss can be propagated back with respect to the controllable parameters. This strategy obviates the need for tedious modeling and analysis of on-chip noise sources, incorporating the noisy behavior of photonic chips naturally during training.
Figure 9: Optical neural network training methods. (a) Normal offline training using a digital platform, e.g., GPU, mimicking the analog hardware behavior by building a physical PIC model. (b) Physical neural network training. It utilizes PICs for forward pass and incorporates a differentiable digital model for backpropagation. (c) In-situ optical neural network training fully on photonic hardware.
@@ III-D2 In-situ Training
The adjoint-variable method was proposed theoretically to implement on-chip backpropagation by interfering with adjoint and forward fields [138]. Recent work has implemented the concept of in situ backpropagation within a triangular MZI mesh [139]. As depicted in Fig. 10, the method involves forward and backward signal propagation, followed by gradient calculation. In step 1 and step 2, the “forward inference” signal $x$ and “backward adjoint” signal $x_{aj}$ are sent forward, respectively. Then the “sum” vector $x-i(x_{aj})^{*}$ is send forward again. The gradient with respect to the phase variables can be finally obtained. However, this implementation requires additional power/phase monitors, faster microcontrollers, and more precise detectors, which increases hardware control complexity and imposes scalability concerns. Additionally, Ohno et.al attempted to design an MRR crossbar array capable of implementing matrix multiplication between the gradient and the transpose of the weight matrix [112], which is a fundamental operation in conventional backpropagation algorithms. While promising, this method remains in the simulation without on-chip demonstrations. Besides, there is a line of works on on-chip learning protocols evaluated on simulation, which will be discussed in Section V-D.
Figure 10: In situ backpropagation concept of adjoint-variable method [139]. The forward, backward, and sum steps of their backpropagation concept are shown to derive the gradient with respect to phase parameters $\theta$.
@@ IV-A1 Fully-Optical accelerator
A fully optical accelerator refers to implementing all operations, including both computation and necessary nonlinear activation functions, entirely within the optical domain. Recent works have demonstrated the integration of optical computation and on-chip nonlinearity [140, 141, 142, 34, 143, 144]. Fully optical accelerators, although promising high bandwidth without the power consumption on electronic interference, still confront significant challenges due to scalability and practical implementation concerns. Firstly, the on-chip nonlinearity is not yet a mature scheme with low energy efficiency or flexibility compared to electronics, as discussed in Section III-C2. Secondly, the significant optical signal loss imposes a substantial requirement for optical power to meet the detection threshold. Additionally, computation errors will accumulate along the optical computing layers, deviating the final outcome significantly from the expected value.
@@ IV-A2 A more practical paradigm: Photonic-electronic Accelerator
Given the challenges associated with practically implementing all-optical accelerators, the photonic-electronic hybrid accelerator emerges as a more feasible and competitive photonic AI solution [145], which is also the key focus in this review. The current hybrid accelerator setup takes advantage of both emerging photonics and mature electronics and builds a system with a tight integration of photonic and electronic integrated circuits. The intensive GEMM operations are executed on optical parts in the analog domain, while the electronic part comprises digital memory for data storage and distribution and necessary digital units for data write/read, flow control, and slight data processing. Combining digital and analog domains results in a mixed-signal setup, therefore, requiring E-O/O-E, D/A, and A/D conversions. Although the above signal conversion brings extra power consumption, it also provides some advantages. For example, the A/D conversion process can be viewed as a denoising step, as it involves discretizing the continuous analog signals into digital representations. This discretization helps filter out noise that is present in the analog signals, avoiding error propagation.
@@ IV-B1 Architecture
The photonic-electronic accelerator can be classified into two types based on its application, which range from task-specific accelerators, such as those for CNN [146, 147, 148], to general-purpose architectures [149, 114]. Despite their varied applications, these accelerators exhibit a similar generic high-level micro-architecture, as shown in Fig. 11. The figure describes the generic architecture of a photonic accelerator in a top-down manner, from the top view of an on-chip accelerator to interaction with peripheral devices, such as off-chip Dynamic Random-Access Memory (DRAM). Within the accelerator, the chip node contains both the photonic and electronic circuits. The photonic part handles tensor multiplication operations, while the electrical component is responsible for flow management. Each tile is constructed with multiple PTCs and a shared local Static Random-Access Memory (SRAM) to store data that can be accessed by different PTCs. The PTC necessitates several digital units to support its working, such as the in-buffer and out-buffer for data storage, ADC, DAC, and the modulator for signal conversion.
@@
Memory. Memory design is indeed a critical aspect of modern accelerator design, as data movement often constitutes the bottleneck of the entire system, a challenge observed not only in traditional electronic processors but also in optical processors [150]. In addition to addressing costly data movement, the ultra-fast characteristic of photonic computing requires specialized memory design to accommodate the requirements for high-speed data access.
@@
Several key design considerations can be taken into account during the memory system design process: (1) Memory Hierarchy: Adopting a memory system with multiple hierarchical levels is a common strategy to reduce memory access costs and meet the speed requirements of the data. Previous studies have organized memory systems from external off-chip DRAM to internal on-chip SRAM [149, 151, 114, 146, 147]. This progression typically involves using global SRAM to drive multiple local SRAMs, with the local SRAM storing frequently accessed data. For example, Lightmatter introduces an architecture, ADEPT [149] as shown in Fig. 11.d, with two separate activation SRAM and weight SRAM, as activation and weights show different access frequencies and patterns. Specifically, weight buffers are utilized to communicate large but slow global SRAM and ultra-fast PTC to achieve parallel weight block programming. (2) Bandwidth matching: Towards efficient data transfer, bandwidth matching is critical to ensure that the memory bandwidth aligns with the requirements of PTCs. One effective approach is to integrate multiple memory channels, as demonstrated in Ref [149, 114]. To meet latency demands and enable efficient transmission of parameter blocks, the large global SRAM block is partitioned into smaller sub-arrays, allowing data to be read at brief intervals to align actual data access latency with specified requirements. (3) Data prefetching: Given the regularity of neural network workloads, future memory access can be anticipated, and double buffering can be employed to hide data loading latency, as utilized in Ref [114]. (4) Data reuse: Optimize the memory system to maximize data reuse and minimize data movement. Weight static is a typical approach in photonic accelerators to explore weight locality and reuse.
@@
Optical Computation. The fundamental computing primitive is a single PTC designed to accelerate tensor operations. The accelerator could integrate multiple PTCs, allowing for task distribution across multiple cores, thus reducing execution time and increasing throughput. These multiple on-chip PTCs can be further organized into tiles, following a modular design principle where each tile shares associated memory and control logic. Different tiles communicate with each other through router logic, using the network-on-chip or optical interconnect [114], as demonstrated in Fig. 11.a.
@@
Digital Electronic Processing Unit. In neural network acceleration, various non-GEMM operations are essential, such as element-wise non-linear operations (e.g., ReLU, GELU, and sigmoid), reduction operations (e.g., softmax and max-pool), batch and layer normalization, and element-wise multiplication and addition (e.g., bias). Digital electronic processing units are employed for these tasks due to their higher efficiency. However, matching the throughput of the photonic part with that of the digital electronics part is crucial to retaining the performance gain due to photonics, especially considering the high throughput of the optical computing part compared to the challenge of electronic processing units operating beyond 2 GHz. Demirkiran et al. [149] propose to equip each PTC with vectorized processing units and duplicate logic units internally to enhance processing speed.
@@ IV-B2 Workload Mapping
This subsection provides an overview of how photonic accelerators perform neural network inference, especially when handling matrix-multiplication-related workloads. The workload needs to be partitioned and scheduled appropriately to map the matrix multiplication on the PIC platform. How to tile and map workloads on the hardware is referred to as mapping.
The first mapping method is specific to convolutional layers, where the mapping honors the sliding window property of convolution. The sliding of the input feature map through the kernel can be viewed as a dot-product with the kernel matrix and the sliding receptive field. Ref [152, 148] explicitly implement the vector dot-product with their photonic circuits. Specifically, in Ref [32], the sliding receptive fields are sent to the photonic accelerator with a time-wavelength interleaving to enable the multiplication with kernel weights in a temporal manner. Ref [153, 146, 147] further cascade multiple sliding receptive fields with the same kernel to form a matrix-vector multiplication. However, this mapping method is specific to convolutional layers and leads to corresponding hardware implementation, limiting its generalizability to handle non-convolutional workloads.
The alternative mapping approach involves transforming the tensor multiplication task into a general matrix multiplication (GEMM) operation between two large matrices. This conversion is straightforward for linear layer workloads. In the case of convolutional layers, the image-to-column (img2col) technique is commonly employed to convert them into GEMM tasks [154, 155]. Matrix tiling is required to partition the input matrices into smaller blocks, and the block-wise multiplication is executed on the photonic accelerator. However, the scheduling of these small blocks, i.e., the order in which these tiles are processed in PTCs, requires careful design, which is sometimes overlooked in recent optical computing studies. Dataflow selection has a significant impact on this scheduling. Currently, many optical computing platforms constrain dataflow selection to weight-stationary dataflow [29, 149]. This is because weight programming is often slow and costly, favoring weight-static mode. In this scheme, block-wise weight matrices are kept in the PTC to maximize their reuse among different input data sizes, and the input data are switched. However, recent advancements in dynamically-operated PTC designs, such as those proposed in Ref [114, 156], have eliminated these restrictions due to advancements in weight programmability, enabling dataflow selection flexibility based on workload requirements. For example, output-stationary dataflow is employed in Ref [114] to support attention matrix multiplication, as the matrices have limited reuse opportunity in this dynamic matrix multiplication scenario.

### v2

@@ III-C2 All-optical nonlinearities
Without electrical components, all-optical approaches implement activation functions via the nonlinear response of materials or devices to optical signals. Several ONNs have been demonstrated using semiconductor optical amplifiers (SOAs) [141], saturable absorbers [142, 143], and techniques exhibiting excitable behavior [144]. All-optical nonlinearities can also be achieved by PCMs, such as the all-optical neuron developed by Feldmann et al [32]. In this work, the PCM, functioning as a waveguide cladding, is used to modulate the pre-synaptic input and govern the resonance state of MRR, thereby controlling the generation of output spikes (Fig. 8.c). Beyond the issues of activation function applicability previously discussed, the implementation of all-optical neurons faces several other challenges. Firstly, even though extra electrical devices are not required, the substantial power required to excite nonlinearities does not offer any power consumption advantages compared to the O-E-O approaches. Additionally, the fabrication and integration of optical components for all-optical nonlinearity, such as PCM, excitable lasers, and optical amplifiers, pose challenges as well. Lastly, since both the control signal and post-synaptic output are optical signals, a potential issue exists in distinguishing the response from the control signals or bias.
@@ III-D Optical Neural Network Training
As an analog computing platform, PICs for AI applications are inherently susceptible to various non-ideal conditions, such as environmental changes, process variation, and limited control precision. These factors can decrease inference accuracy and potentially degrade the signal-to-noise ratio (SNR) of computations. To mitigate the decrease in accuracy, it is crucial to train ONNs with careful consideration of non-idealities that may occur during inference, making it a “circuit-aware” approach will enhance its noise resilience. The current ONN training can be classified into two categories: hardware-aware ex-situ training, which involves training with the help of digital computers, and on-chip in-situ training.
@@ III-D1 Hardware-aware ex-situ Training
Hardware-aware ex-situ training offloads the training process to digital computers and utilizes various hardware-aware training techniques to capture PIC behavior as precisely as possible during training. One commonly used technique is noise-aware training [145, 146], as shown in Fig. 9.a. This approach involves modeling the behavior of PICs while considering various non-ideal effects. Subsequently, the PIC model is injected into the training to reduce the gap between training and real inference. Gaussian noise is commonly used to model various noise sources in photonic systems. For instance, shot noise and thermal noise are modeled using a Gaussian distribution by measuring on-chip photonic multiplication results and fitting their distribution [132]. In addition to injecting non-idealities into the training, some work explicitly models the transfer matrix of photonic neurons and embeds them in the forward computation pass during training [130, 131]. This injection is crucial because it introduces a unique nonlinearity term that ONNs need to be aware of in order to accurately capture the behavior of photonic neurons. Moreover, in Ref [126], the noisy transfer matrix is derived and explicitly injected in the training under some noise assumptions. However, this approach encounters two main challenges. Firstly, accurately representing all on-chip non-idealities poses a significant difficulty, and environmental fluctuations can further affect inference accuracy. Second, the computational overhead required to model the physical system accurately can be exceedingly high or even prohibitive, making the training very slow and costly.
Aware of these challenges, some works advocate for training optical neural networks directly with non-ideal physical responses, called physical network training, as shown in Fig. 9.b. The forward pass is executed on PICs and the loss signal $\mathcal{L}$ is obtained by comparing the physical and intended outputs. However, a challenge arises during the backward pass due to the undifferentiable nature of the physical response. To address this challenge, the idea of adopting a differentiable PIC surrogate model in digital domain has been proposed [89, 147, 148]. With the differentiable model, the gradient of the loss can be propagated back with respect to the controllable parameters. This strategy obviates the need for tedious modeling and analysis of on-chip noise sources, incorporating the noisy behavior of photonic chips naturally during training.
Figure 9: Optical neural network training methods. (a) Offline training on a digital platform, such as a GPU, mimics the behavior of analog hardware by building a physical model of PICs. (b) Physical neural network training, which utilizes PICs for forward pass and incorporates a differentiable digital model for backpropagation. (c) In-situ optical neural network training is performed entirely on photonic hardware.
@@ III-D2 In-situ Training
Although in-situ training is straightforward and ideal, it is challenging to implement directly on the optical computing platform. Considering an optical neural network layer $l$, optical components execute the linear projection $W^{l}(\mathbf{\Phi}^{l})$ complemented by a digital or analog nonlinear transformation $f^{l}$. With the input $\mathbf{x}^{l}$, the forward manner can be defined as,
[…]
Here, $\mathbf{\Phi}^{l}$ represents the device configurations, which are the device control variables. For the backward pass, assuming we can access the gradient of loss over $\mathbf{x}^{l+1}$, ${\partial\mathcal{L}}/{\partial\mathbf{x}^{l+1}}$, we need to obtain the gradient over input $\mathbf{x}^{l}$ and device configurations $\mathbf{\Phi}^{l}$. The first step is to obtain the gradient over $\mathbf{y}^{l}$, which could be prohibitive if the activation function is undifferentiable without analytical formulations, especially for customized analog activations. After determining the gradient ${\partial\mathcal{L}}/{\partial\mathbf{y}^{l}}$, the gradients over inputs and device configurations are defined as $W^{l}(\mathbf{\Phi}^{l})^{T}\frac{\partial\mathcal{L}}{\partial\mathbf{y}^{l}}$ and $\frac{\partial\mathcal{L}}{\partial\mathbf{y}^{l}}\mathbf{x}^{T}\frac{\partial\mathbf{W}^{l}(\mathbf{\Phi}^{l})}{\mathbf{\Phi}^{l}}$. The challenges of obtaining the above two items stem from several aspects. First, it requires bidirectional input support to access the transpose of $W^{l}(\mathbf{\Phi}^{l})$. Second, one may argue that another optical processor can be used to implement matrix multiplication in the backward pass, thereby avoiding the need for bidirectional support. However, noise and precision limitations cause the gradient computation to deviate from the desired matrix multiplication, thus impeding convergence. Third, obtaining the analytical gradient over the real device configurations, ${\partial\mathbf{W}^{l}(\mathbf{\Phi}^{l})}/{\mathbf{\Phi}^{l}}$, can be very complex and prohibitive, as demonstrated in the case of an MZI-ONN [149].
An adjoint-variable method was proposed theoretically to implement on-chip backpropagation by interfering with adjoint and forward fields [150]. Recent work has implemented the concept of in-situ backpropagation within a triangular MZI mesh [151]. As depicted in Fig. 10, the method involves forward and backward signal propagation, followed by gradient calculation. In step 1 and step 2, the “forward inference” signal $x$ and “backward adjoint” signal $x_{aj}$ are sent forward, respectively. Then the “sum” vector $x-i(x_{aj})^{*}$ is send forward again. The gradient with respect to the phase variables can be finally obtained. However, this implementation requires additional power/phase monitors, faster microcontrollers, and more precise detectors, which increases hardware control complexity and imposes scalability concerns. Additionally, Ohno et al. attempted to design an MRR crossbar array capable of implementing matrix multiplication between the gradient and the transpose of the weight matrix [123], which is a fundamental operation in conventional backpropagation algorithms. While promising, this method has yet to be demonstrated on-chip. Besides, there is a line of works on on-chip learning protocols evaluated in simulation, which will be discussed in Section V-D.
Figure 10: In-situ backpropagation concept of adjoint-variable method [151]. The forward, backward, and sum steps of their backpropagation concept are shown to derive the gradient with respect to phase parameters $\theta$.
@@ IV-A1 Fully-optical accelerator
A fully optical accelerator refers to implementing all operations, including both computation and necessary nonlinear activation functions, entirely within the optical domain. Recent works have demonstrated the integration of optical computation and on-chip nonlinearity [152, 153, 32, 154]. Fully optical accelerators, although promising high bandwidth without the power consumption of E-O interference, still confront significant challenges due to scalability and practical implementation concerns. Firstly, the on-chip nonlinearity is not yet a mature scheme with low energy efficiency or flexibility compared to electronics, as discussed in Section III-C2. Secondly, the significant loss imposes a substantial requirement for optical power to meet the detection threshold. Additionally, computation errors will accumulate along the optical computing layers, deviating the final outcome significantly from the expected value.
@@ IV-A2 A more practical paradigm: photonic-electronic accelerator
Given the challenges associated with practically implementing all-optical accelerators, the photonic-electronic hybrid accelerator emerges as a more feasible and competitive photonic AI solution [155], which is also the key focus in this section. The current hybrid accelerator setup takes advantage of both emerging photonics and mature electronics and builds a system with a tight integration of photonic and electronic integrated circuits. The intensive tensor operations are executed on optical parts in the analog domain, while the electronic segment includes digital memory for data storage and distribution, as well as essential units for data writing/reading, flow control, and minor data processing. Combining digital and analog domains results in a mixed-signal setup, therefore, requiring E-O/O-E, D/A, and A/D conversions. Although the conversion processes incur additional power consumption and delay, they also offer some advantages. For example, the A/D conversion process can be viewed as a denoising step, as it involves discretizing the continuous analog signals into digital representations. This discretization helps filter out noise that is present in the analog signals, avoiding error propagation.
@@ IV-B1 Architecture
The photonic-electronic accelerator can be classified into two types based on its application, which range from task-specific accelerators, such as those for CNNs [156, 157, 158], to general-purpose architectures [159, 126]. Despite their varied applications, these accelerators exhibit a similar generic high-level micro-architecture, as shown in Fig. 11. The figure describes the generic architecture in a top-down manner, from the top view of an on-chip accelerator to interaction with peripheral devices, such as off-chip dynamic random-access memory (DRAM). Within the accelerator, the chip node contains both the photonic and electronic circuits. The photonic part handles tensor operations, while the electrical component is responsible for flow management. Each tile is constructed with multiple PTCs and a shared local static random-access memory (SRAM) to store data that can be accessed by different PTCs. The PTC necessitates several digital units to support its operation, such as the in-buffer and out-buffer for data storage, ADC, DAC, and the modulator for signal conversion.
@@
Memory Memory design is a critical aspect of modern accelerator design, given that data movement frequently constitutes the bottleneck of the entire system. This challenge is observed not only in traditional electronic processors but also in optical processors. [160]. In addition to addressing costly data movement, the ultra-fast characteristic of photonic computing requires specialized memory design to accommodate the requirements for high-speed data access.
@@
Several key design considerations can be taken into account during the memory system design process: (1) Memory hierarchy: Adopting a memory system with multiple hierarchical levels is a common strategy to reduce memory access costs and meet the speed requirements. Previous studies have organized memory systems from external off-chip DRAM to internal on-chip SRAM [159, 161, 126, 156, 157]. This progression typically involves using global SRAM to drive multiple local SRAMs that store frequently accessed data. For example, Lightmatter has introduced an architecture denominated ADEPT as shown in Fig. 11.d [159], featuring separate SRAMs for activations and weights to accommodate their differing access frequencies and patterns. Specifically, weight buffers are utilized to facilitate communication between the large but slow global SRAMs and the ultra-fast PTCs, enabling parallel programming of weight blocks. (2) Bandwidth matching: Towards efficient data transfer, bandwidth matching is critical to ensure that the memory bandwidth aligns with the requirements of PTCs. One effective approach is to integrate multiple memory channels, as demonstrated in Ref [159, 126]. To meet latency demands and enable efficient transmission of parameter blocks, the large global SRAM block is partitioned into smaller sub-arrays. This allows data to be read at brief intervals, aligning actual data access latency with specified requirements. (3) Data prefetching: Given the regularity of neural network workloads, future memory access can be anticipated, allowing for the use of double buffering to hide data loading latency, as demonstrated in Ref [126]. (4) Data reuse: A common approach in photonic accelerators, known as ’weight static’, explores weight locality to enhance reuse and minimize data movement.
@@
Optical computation The fundamental computing primitive is a single PTC designed to accelerate tensor operations. The accelerator could integrate multiple PTCs, allowing for task distribution across multiple cores, thus reducing execution time and increasing throughput. These multiple on-chip PTCs can be further organized into tiles, following a modular design principle where each tile shares associated memory and control logic. Different tiles communicate with each other through router logic, using network-on-chip or optical interconnect, as demonstrated in Fig. 11.a.
@@
Digital electronic processing unit In neural network acceleration, various non-GEMM operations are essential, such as element-wise non-linear operations (e.g., ReLU, GELU, and sigmoid), reduction operations (e.g., softmax and max-pool), batch and layer normalization, and element-wise operations (e.g., bias). Digital electronic processing units are employed for these tasks due to their higher efficiency. However, to retain the performance benefits of photonics, it is crucial to match the high throughput of photonic part, especially considering the challenges faced by electronic processing units operating beyond 2 GHz. Demirkiran et al. proposed to equip each PTC with vectorized processing units and duplicate logic units internally to enhance processing speed [159].
@@ IV-B2 Workload Mapping
This subsection provides an overview of how photonic accelerators perform neural network inference, especially when handling matrix-multiplication-related workloads. The workload needs to be partitioned and scheduled appropriately to map the matrix multiplication to PICs.
The first mapping method is specific to convolutional layers, where the mapping honors the sliding window property of convolution. The sliding of the input feature map through the kernel can be viewed as a series of dot products between the kernel matrix and the sliding receptive fields. In Ref [30], the sliding receptive fields are sent to the photonic accelerator with a time-wavelength interleaving to enable the multiplication with kernel weights in a temporal manner. Ref [156, 157] further cascade multiple sliding receptive fields with the same kernel to perform an MVM operation. However, this mapping method is specific to convolutional layers, which limits its generalizability across non-convolutional workloads.
The alternative mapping approach involves transforming the tensor multiplication task into a GEMM operation between two large matrices. This conversion is straightforward for linear layer workloads. In the case of convolutional layers, the image-to-column (img2col) technique is commonly employed to convert them into MVM tasks [162, 163]. Matrix tiling is required to partition the input matrices into smaller blocks, and the block-wise multiplication is executed on the photonic accelerator. However, the scheduling of these small blocks, i.e., the order in which these tiles are processed in PTCs, requires careful design, which is sometimes overlooked in recent optical computing studies. Dataflow selection has a significant impact on this scheduling. Currently, many optical computing platforms constrain dataflow selection to weight-stationary dataflow [28, 159]. This is because weight programming is often slow and costly, favoring the weight-static mode. In this scheme, block-wise weight matrices are kept in the PTC to maximize their reuse among different input data sizes. However, recent advancements in dynamically-operated PTC designs, such as those proposed in Ref [126, 164], have eliminated these restrictions due to advancements in weight programmability, enabling flexible dataflow selection based on workload requirements. For example, output-stationary dataflow is employed in Ref [126] to support attention matrix multiplication, as the matrices have limited reuse opportunity in this dynamic matrix multiplication scenario.